OpenStudy (anonymous):

During a hike, Vijay saw a total of 13 birds and butterflies. Knowing that birds have 2 wings and butterflies have 4 wings, he counted 46 wings. How many birds did he see? A. 1 B. 3 C. 6 D. 8

OpenStudy (anonymous):

x=birds y=butterflies x+y=13 2x+4y=46 so... x=13-y 2(13-y)+4y=46 26-2y+4y=46 2y=20 y=10 x=13-10 x=3 3 birds